# Designing quantum chemistry algorithms with just-in-time compilation

### v5

@@ 5 Conclusion
We optimized the low–angular-momentum electron repulsion integral algorithm in GPU4PySCF v1.4 using just-in-time (JIT) compilation and introduced a novel high–angular-momentum integral method. The core CUDA implementation for both algorithms spans only 1,000 lines of code in total. JIT compilation not only accelerates the evaluation of Gaussian-type-orbital integrals but also provides developers with greater flexibility. Our streamlined implementation delivers a 2-4x performance improvement over GPU4PySCF v1.4 and is significantly easier to maintain. Additionally, it supports single- and double-precision arithmetic, accepts multiple density matrices as input, and integrates seamlessly with the existing GPU4PySCF framework.
@@ 6 Future Work and Limitations
Supporting more functionalities. A practical quantum‐chemistry workflow requires more than the acceleration of Coulomb and exchange kernels. The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post Hartree-Fock methods, and a wide array of molecular properties are all essential. By integrating JoltQC with GPU4PySCF, users can build sophisticated simulation pipelines, and any of these routines can be further accelerated on demand using JIT compilation.
Mixed precision algorithms. This paper focuses on the performance of the double precision and single precision separately. Thanks to the Schwartz inequality, the computational load of the integrals can be naturally split into FP64 calculations and FP32 calculations without loss of accuracy theoretically. Other research efforts [17, 20] show that 90% of the computations can be in FP32. A naive mixed precision strategy can be easily proposed based on this work. One should notice, even though the value of integral can be represented in FP32, not every underlying arithmetic operation for computing the integral can necessarily be computed with FP32. More effort is needed to rigorously implement a strategy without error accumulation.
Screening techniques. The intrinsic scalings of the Coulomb and exchange terms are different. The number of significant Coulomb terms increases quadratically, while the number of significant exchange terms increases linearly with respect to the system size. It is more efficient to evaluate the Coulomb and exchange terms separately. In some quantum chemistry packages such as TeraChem, the Coulomb terms are efficiently evaluated with J-engine [32]. Besides the Schwartz inequality, which gives a rigorous upper bound, other estimates can further eliminate the unnecessary calculations. For example, the MBIE [14], QQR [18], and CSAM [27]. MBIE provides a rigorous upperbound, but it is expensive to calcuate. QQR and CSAM take advantage of distance dependence in their estimates, hence improving the performance for large systems. But those estimates are less rigorous.

### v6

@@ 5 Conclusion
We have demonstrated that just-in-time compilation is a powerful paradigm for accelerating electron repulsion integral evaluation on GPUs. By making angular momentum, orbital contraction pattern, and primitive counts compile-time constants, the JIT compiler can aggressively unroll loops, eliminate branches, and optimize register allocation — yielding a 2× speedup for small basis sets (6-31G*) and up to 4× for large basis sets (def2-TZVPP) over GPU4PySCF v1.4 on an A100-80G GPU. In single precision on an A10-24G, JoltQC is approximately three times faster than TeraChem v1.9 for the JK kernels.
Beyond performance, JIT compilation fundamentally changes the development model: the core NVRTC implementation spans only 1,000 lines for both algorithms, compared to $\sim$20,000 lines in GPU4PySCF v1.4. This compactness is not incidental — it is a direct consequence of delegating specialization to the compiler rather than encoding it by hand. The resulting code is easier to maintain, extend, and adapt to new hardware. We expect JIT compilation to become a standard tool in the quantum-chemistry software stack as GPU architectures continue to diversify.
We release JoltQC as an open-source library to serve as both a practical tool and a foundation for applying JIT compilation more broadly in quantum chemistry. JoltQC supports single- and double-precision arithmetic, accepts multiple density matrices, and integrates seamlessly with GPU4PySCF.
@@ 6 Future Work and Limitations
Supporting more functionalities. A practical quantum‐chemistry workflow requires more than the acceleration of Coulomb and exchange kernels. The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post-Hartree-Fock methods, and a wide array of molecular properties are all essential. By integrating JoltQC with GPU4PySCF, users can build sophisticated simulation pipelines, and any of these routines can be further accelerated on demand using JIT compilation.
Mixed precision algorithms. This paper focuses on the performance of the double precision and single precision separately. Thanks to the Schwartz inequality, the computational load of the integrals can be naturally split into FP64 calculations and FP32 calculations without loss of accuracy theoretically. Other research efforts [18, 21] show that 90% of the computations can be in FP32. A straightforward mixed-precision strategy follows directly from the Schwarz inequality: shell quartets whose integral upper bound exceeds the FP32 representable range are computed in FP64, while the remaining 90% are computed in FP32. JoltQC’s JIT framework makes this trivial to implement — the precision is already a template parameter (Table I), so both FP32 and FP64 kernels are compiled and dispatched based on the screening estimate. A key subtlety is that even when the final integral value lies within FP32 representable range, the intermediate arithmetic — particularly the recursive evaluation of Rys polynomials and the accumulation of primitive contributions — may involve catastrophic cancellation or large intermediate values that require FP64. A rigorous mixed-precision strategy must therefore identify which computational steps can safely use FP32 and which must remain in FP64 to avoid error accumulation. This is a nontrivial compiler-level analysis that we leave to future work.
Screening techniques. The intrinsic scalings of the Coulomb and exchange terms are different. The number of significant Coulomb terms increases quadratically, while the number of significant exchange terms increases linearly with respect to the system size. It is more efficient to evaluate the Coulomb and exchange terms separately. In some quantum chemistry packages such as TeraChem, the Coulomb terms are efficiently evaluated with J-engine [33]. Besides the Schwartz inequality, which gives a rigorous upper bound, other estimates can further eliminate the unnecessary calculations. For example, the MBIE [15], QQR [19], and CSAM [28]. MBIE provides a rigorous upperbound, but it is expensive to calculate. QQR and CSAM exploit distance dependence in their estimates, offering tighter bounds for large systems at the cost of reduced rigor compared to the Schwarz inequality. Integrating these screening strategies into JoltQC is straightforward: because the screening is performed on the host before kernel dispatch, it is independent of the JIT-compiled kernels and can be swapped without recompilation.